Selling bonds.  Lunar Vacations needs to raise $6,200,000 for its new project​ (a golf course on the​ moon). Astro Investment Bank will sell the bond for a commission of 2.3%.

The market yield is currently 7.1% on​ twenty-year semiannual bonds. If Lunar wants to issue a 6.2% semiannual coupon​ bond, how many bonds will it need to sell to raise the

$6,200,000​? Assume that all bonds are issued at a par value of $1,000.

How many bonds will Lunar need to sell to raise the $6,200,000​?
Answer #1

First we find the price of the bond

Using financial calculator
Input: FV= 1000, I/Y=7.1/2

N=20*2=40

PMT=6.2%*1000/2= 31

Solve for PV as -904.64

Commission on above price= 2.3%*904.64 = 20.81

hence net proceeds to the company per bond = 904.64-20.81 = 883.84

Number of bonds to be sold = 6200000/883.84 = 7014.84

rounded off to 7015 bonds